NAMMA Banner

NPOL Radar report

September 8th, 2006

08 Sept 06
Status; Green
Radar working well.

Heavy convection associated with wave passed over the site yesterday evening.
There was evidence this morning of circulation indicating the possible development of a vortex.
Attached is a radar image from 0945 UTC Today

NPOL 08/31/06

September 1st, 2006

Date: 08/31/06


Instrument Status:  Green


Instrument Performance:

NPOL continues to run well.  We had a short issue with a bad connection in the slip ring for a time early on the 31st, but it was resolved.  NPOL ran well for the precipitation event later that day.


Science Observations:

Around 0430 UTC NPOL began to detect a mesoscale convective system (MCS) in the NE quadrant during surveillance scans.  The line appered to be in a near mature state as it entered the NPOL scanning area.  The line progressed in a southwest direction, around 60 km/hr.  It increased in strength and size as it approached the NPOL and eventually developed a large stratiform rain region.  At 0700 UTC, the leading convective line appeared as if it would bow and portions had reflectivity values greater than 50 dBZ. At 0730 UTC, the intensity of the leading convective line decreased somewhat and there were small gaps in the reflectivity of the trailing stratiform rain region. However, the MCS appeared to go through a regeneration phase as a new line of cells developed along the southeast flank of the MCS. The system crossed the radar shortly after 0800 UTC, with the gust front passing at 0804 UTC and rain commencing shortly thereafter. High winds were associated with the passing of the squall line with maximum wind gusts reaching 46 mph. The system continued to propagate out over the ocean although it began to weaken again with decreased reflectivity values in the leading line. While over land, the MCS crossed nearly all of the rain gauge sites, including the dense rain gauge network.  By 1600 or 1700 UTC the system had completely departed from the NPOL scanning area.

A. Newman and A. Reynolds

NPOL 08/26/06

August 26th, 2006

Date: 08/26/2006

Instrument Status: Green

Instrument Performance:

NPOL continues to run well.  We did shutdown for generator maintanance for a 1.5 hours yesterday and for a few scans to update the bias offset in the ZDR field.

Communications remains the biggest problem.  We don’t expect an internet connection at the radar until next week at the earliest.  Email connect at a local internet cafe remains our only connection except by cell phone.
Science Observations:

We had our first day of activity since operations started on Monday.  Most of the convection was scatter and unorganized.  The cells tended to move to the N and rotate to the NW at our maximum range (150 km).  The intensity of the storm was week with maximum reflectivity rarely exceeding 40 dBZ and echo top heights ranged between 7-10 km.

We coordinated with the DC-8 between 1645-1730 UTC as the plane flew a small line to our south.  Below is summary of the observations from NPOL:

1645 UTC: NASA436 made contact with NPOL.  They worked a small line that was located about 30 km to the SSW of NPOL.  It was orientated from WNW to SSE.  The length of the line is approximately 75 km.  Maximum reflectivity was about 43 dBZ with echo top heights not exceeding 8 km.

1700 UTC:  The line continues to move to the NNE at about 10 km/h.  The DC-8 is in a holding pattern waiting clearance.  They are flying near a smaller line to the east of the main line, which is located about 35 km SE of NPOL.

1715 UTC: The DC-8 has been cleared to work the line to the SSW of NPOL.  The line continues to move north and maintain the vertical structure and intensity as observed in the previous scans.

1731 UTC: The DC-8 is finishing up the last segment and moving out for the flight back to CV.

1815 UTC: The line has moved over NPOL and dissipated except for a few isolated cells.

Note: the Radio communications worked very well during the mission.

P. Kucera

NPOL 8/21/06

August 21st, 2006

Date: 08/21/2006

Instrument Status: Green

Instrument Performance:

NPOL is now operational as of about 1415 UTC today.  We will be operating 24/7 unless we encounter any problems.

Science Observations:

No significant weather today.

P. Kucera


NPOL Daily Summar

August 20th, 2006

Date: Sunday 08/20/2006
Instrument Status: Yellow
Instrument Performance:NPOL is now operational. However, we have a limited fuel supply for the generator until Monday. We were calibrated yesterday afternoon and observed the squall line as it transitioned over the ocean for a couple of hours. We plan to be operational 24/7 starting around 9 AM UTC tomorrow.

Daily Summary Template

July 18th, 2006

Date: 07/15/2006

Instrument Status: Red, Yellow, Green

Quicklook(s): Sample Quicklook

Instrument Performance:

How did the instrument perform during the flight?

Science Observations:

A brief report on anything interesting.

Current Status Report Template

July 12th, 2006

Date: 07/15/2006
Time: 00:00 (UTC)
Current Status:

A brief report on the status of the instrument at the current time.


Have you used our data? Register for updates